How do you solve:
Chris Siragusa takes out a 7% simple interest loan today which will be repaid 15 months from now with a payoff amount of $815.63. What amount is Chris borrowing?
Thank you very much!

Answer by mananth(16949)

The basic formula for Interest is
I = p*r*t
let loan amount be $x
at 7% for 1.25 years
interest = 0.075x* 1.25
...
0.07*1.25*x +x =815.63
1.0875x = 815.63
x = 815.63/1.0875
x = $750
...
m.ananth@hotmail.ca